### v1

@@ 4.1.1 Participants
Two healthy men, aged 39 and 30 years, were recruited for this study. Both participants had no history of neurological or musculoskeletal disorders that could affect motor neuron functionality. The small subject sample was due to the relatively complex implant procedure, which included three micro-electrode arrays per subject, with ultrasound-guided insertion (see below). All experimental procedures adhered to the ethical guidelines set by Imperial College London. The studies were performed according to the Declaration of Helsinki, with an informed consent form signed by all participants before each experiment (ICREC Project ID 19IC5640 for the intramuscular recordings; JRCO Project ID 18IC4685 for the MRI acquisition).
@@ 4.1.2 Experiments
HD-iEMG signals were recorded from forearm muscles using multi-channel intramuscular electrodes designed for acute recordings [13, 12] (Fig. 6 a). Additionally, three surface electrode grids (IED 4mm) with 64 recording sites each (Fig. 6 c) were positioned over the targeted muscles to concurrently acquire HD-sEMG signals. This configuration enabled concurrent measurement of the electrical activity both within the muscle and at the skin surface. The EMG signals were acquired using a multi-channel amplifier (OT-Bioelettronica, Torino, Italy). All EMG data were sampled at 10,240 Hz, and digitised with 16-bit resolution. The HD-EMG signals were bandpass filtered (Butterworth, order 5) between 100-4400 Hz (HD-iEMG) and 20-500Hz (HD-sEMG). Recordings were made in bipolar configuration with a reference electrode placed on the wrist.
The participants performed force-tracking tasks following trapezoidal force trajectories at 15 % of their MVC (Fig. 3 b). These tasks were executed individually for each finger (thumb, index, middle, ring, and little finger) in flexion and extension. Additionally, forces were exerted by combinations of fingers (thumb and index; index and middle; middle and ring; index with middle and ring; thumb with index, middle, ring, and little finger) in flexion and extension, following the same trapezoidal profile. During all tasks, participants maintained isometric contractions focusing on the specified movements. The duration of the constant-force part of the contractions was 20 seconds. Two repetitions were recorded for each task.
@@
We classified the different tasks with a series of analyses. First, we attempted the classification of 16 classes that included all individual fingers in flexion and extension, and three combinations of two fingers (T-F, I-F, M-F, R-F, L-F, IM-F, MR-F, IMR-F, T-E, I-E, M-E, R-E, L-E, IM-E, MR-E, IMR-E). Then, we focused on the tasks mostly associated to the targeted muscles, which resulted in 12 classes: I-F, M-F, R-F, IM-F, MR-F, IMR-F, I-E, M-E, R-E, IM-E, MR-E, IMR-E. Finally, as a proof of concept of a specific application of the mapping system - the control of a prosthetic hand - we selected tasks that are functionally important for amputees and challenging to decode with state-of-the-art methods [15, 33]. This group of tasks included power grip (TIMRL-F, TIMRL-E), pinch (TI-F, TI-E), and individual index/thumb finger activation (T-F, T-E, I-F, I-E), for a total of eight classes. All tasks were performed isometrically, with the corresponding fingers activated simultaneously. However, the power grip class was excluded for subject S1 due to the inability to correctly perform the coordinated movements of the thumb, index, middle, ring, and little fingers. Specifically, S1 exhibited oscillations in the force applied by the little finger, with fluctuations of amplitude of around 7 %MVC. These force variations led to inconsistent activation of the motor units, which were subsequently removed by the automatic cleaning process. As a result, the last analysis resulted in six classes for S1 (T-F, I-F, TI-F, T-E, I-E, TI-E) and eight for S2 (T-F, I-F, TI-F, TIMRL-F, T-E, I-E, TI-E, TIMRL-E).
@@ 4.1.3 Instrumented platform
A finger force measurement platform was designed to be adaptable to different hand sizes, adjustable in wrist angle, and equipped with high-resolution and cost-effective load cells to capture force signals. A 10-mm polycarbonate board was used for the platform body because of its light weight. Precision laser-cutting of load cell rails was executed using CAD software. The platform features ten 10-kg TAL220 load cells (two per finger, one for flexion, and one for extension), selected based on the maximum observed male exertion forces of 7-8 kg [34, 35]. The load cells are mounted on adjustable rails with rolling pins and connected to 3D-printed mounts to accommodate varying finger thicknesses. The EMG signals and the force data were acquired simultaneously by the Quattrocento amplifier. Force signals were amplified using Forza pre-amplifiers linked to the amplifier. A 3D-printed holder housed the 10 pre-amplifiers, optimising the setup for accurate force measurement across various hand sizes.
@@ 4.2 Classification
We examined the accuracy in classifying intended motor tasks into predefined categories by comparing two approaches: one that used decomposed motor unit activity (i.e., the timings of motor neuron discharge events) and another that used global features derived from the interference signal (e.g., signal amplitude).
@@
To classify tasks using motor neuron behaviour, the HD-iEMG signals were decomposed into the contributing motor units. The filters associated with each source (motor unit) were aggregated into a comprehensive dictionary of filters for each task. Figure 5 panel a provides a schematic representation of this filter bank. In the diagram, each colour corresponds to a different class label, indicating the specific task being represented. The width of each bar reflects the number of filters associated with that task, illustrating how many motor unit filters contribute to the classification of each specific task. The zoomed-in section within the magenta box offers a closer look at a representative filter from the pooled filter bank. This filter visualises the amplitude of the motor unit across different channels, providing insights into how the neural activity is distributed spatially across the electrode array. Conversely, for global EMG classification, the root mean square of the EMG signals served as the feature for a C-Support Vector Machine classifier (Fig. 5 b). To mimic a real-life mapping system, the calibration phase was conducted using the first recorded repetition, and the testing phase was performed using a repetition recorded later in the day.
4.2.1 Finger task classification based on motor unit behaviour
@@
Motor unit spike trains were extracted through blind source separation of HD-iEMG recordings obtained during isometric finger tasks at 15 %MVC (see Appendix A for more details on EMG decomposition). The tasks included individual finger flexion and extension as well as multi-finger combinations performed isometrically. The decomposition process produced task-specific sets of motor unit filters, which were pooled to create a comprehensive bank of motor unit action potential filters $\boldsymbol{\mathbb{F}}_{i}\in\mathbb{R}^{N\times M}$, where $N$ represents the total number of extracted motor unit filters and $M$ the number of EMG channels. Each motor unit filter $\boldsymbol{\mathbb{F}}_{i}$ in this set was labelled according to the specific finger movement task it was derived from (e.g., thumb flexion corresponding to class 0).
@@
These reference motor unit filters from the 15 %MVC contraction were then applied to the multi-channel iEMG signals $\mathbf{E}j\in\mathbb{R}^{M\times D_{R}}$, where $D_{R}$ is the duration of the recording, recorded during a separate 15 % MVC trial involving all fingers performing both flexion and extension movements. This produced the corresponding discharge timings, represented as IPTs $\mathbf{I}{ij}\in\mathbb{R}^{N\times D_{R}}$ for each motor unit $i$ and task $j$. Each IPT $\mathbf{I}_{ij}$ was clustered to separate the peaks (i.e., the source) from the noise components. Only motor units with robust clustering quality (silhouette value >0.9), physiological discharge periodicity (coefficient of variation <35%), and firing rates within the physiological range for isometric contractions (mean discharge rate <30 Hz) were retained for further processing. The discharge timings of these units were temporally segmented into 100-ms time bins with 50% overlap. This window was selected to be long enough for reliable classification while remaining short enough to minimize delay in real-time control applications. The windowing resulted in a binary matrix $\mathbf{B}\in\mathbb{R}^{N\times K}$, where $K$ represents the number of bins. For each time bin $k$, the corresponding class label was assigned according to the filter $\boldsymbol{\mathbb{F}}_{i}$ that produced the source, creating a final label vector $\mathbf{L}\in\mathbb{R}^{K}$. If no predictions were made for a given bin (i.e., no motor units were active), it was labeled as N/A. Finally, for each time bin, the class with the majority of votes was assigned to that bin. If there were no predictions overall for a bin, it was labeled as N/A and counted as a misclassification in the final accuracy calculation. The method, therefore, relies on unique motor units being consistently active throughout the entire task. Figure 5 a shows a schematic of the method.
@@ 4.2.2 Finger classification based on global EMG
For myoelectric decoding of dexterous hand gestures, a linear SVM classifier with an L2 regularisation penalty factor of 0.1 was employed (Fig. 5 b). The SVM classifier is a widely used method in EMG pattern recognition [36, 37, 38]. The HD-EMG signals were bandpass filtered (Butterworth, order 5) between 100-4400 Hz (HD-iEMG) and 20-500 Hz (HD-sEMG) to remove noise. Global normalisation factors based on the maximum absolute value across the pooled training set were derived and applied to the test set for consistent scaling. Outlier channels with baseline noise exceeding 3 standard deviations from the training mean were removed from all training and testing data. The pre-processed EMG was segmented into 100-ms intervals, with 50% overlap using a sliding window, preparing the bins similarly to how they were organised for the classification based on motor units. The r.m.s. values of the EMG were extracted as the input features. This was done in the same way for both surface and intramuscular high-density recordings. Moreover, to simulate a lower-density iEMG electrode configuration, we selected the first and last channel for each electrode, resulting in an IED of 2 cm (Fig. 6 b). The difference between the two electrodes was computed for each array, resulting in 3 feature channels. This configuration was chosen to emulate structures like the implantable myoelectric sensors (IMES) electrodes, currently being tested to control multifunction hand prostheses [39, 40, 41]. For the surface electromyography (sEMG), we selected eight channels in the central column of the electrode grid, mimicking the configurations found in commercial low-density systems such as the MYOband (Fig. 6 d).
@@ 5 Acknowledgments
A.G. was supported by UK Research and Innovation [UKRI Centre for Doctoral Training in AI for Healthcare grant number EP/S023283/1] and by Huawei Technologies Research & Development (UK) Limited. J.I.P. was supported by project ECHOES (ERC Starting Grant 101077693), by a Consolidación Investigadora grant (CNS2022-135366) funded by MCIN/AEI/10.13039/ 501100011033 and UE’s NextGenerationEU/PRTR funds. S.M. was supported by Chalmers Life Science Engineering Area of Advance. A.K. was supported by NISNEM (EPSRC EP/T020970/1). D.F. was supported by NaturalBionicS (ERC Synergy 810346).
@@ Appendix A Decomposition
The EMG signals were decoded into the discharges of the corresponding motor units via a blind source separation algorithm [42].
Multi-channel EMG signals represent combined mixtures of motor neuron spike trains. Each recorded signal, i.e., each EMG channel, is the summation of the convolution of unknown finite impulse response filters, corresponding to a motor unit action potentials, with their respective sources, i.e., the sequences of motor neuron discharge timings. To make the deconvolution (i.e., EMG decomposition) possible, this combined mixture can be converted into a linear instantaneous mixture by extending the sources to include the original $n$ sources and their delayed versions, with delays ranging from 1 to the filter length $L$. Similarly, the $m$ original observations are extended by a factor $R$ to ensure there are more observations than sources [43].
For each time sample $(k)$, indicating the sources, observations and noise, respectively, with ${s}(k)=\left[s_{1}(k),s_{2}(k),\ldots,s_{n}(k)\right]^{T},\underset{=}{x}(k)=\left[x_{1}(k),x_{2}(k),\ldots,x_{m}(k)\right]^{T}$ and ${n}$ $(k)=\left(n_{1}(k),n_{2}(k),\ldots,n_{n}(k)\right)^{T}$, where $T$ is the transpose operator, the extended model is as follows:
${\widetilde{x}}(k)=[\underset{=}{\widetilde{H}}\widetilde{s}(k)+\underset{=}{\widetilde{n}}(k)]\quad\text{for}\quad k=1,\ldots,D_{R}$ (1)
where $D_{R}$ is the duration of the recording in samples, and with the extended sources, observations and noise as $\widetilde{s}(k)=\left[\widetilde{s_{1}}(k),\widetilde{s_{2}}(k),\ldots,\widetilde{s_{n}}\right.$ $\left.\widetilde{s_{n}}(k)\right]^{T},\widetilde{x}(k)=\left[\widetilde{x_{1}}\widetilde{x_{1}}(k),\widetilde{x_{2}}(k),\ldots,\widetilde{x_{m}}(k)\right]^{T}$ and $\widetilde{n}(k)=\left[\widetilde{n_{1}}(k),\widetilde{n_{2}}(k),\ldots,\widetilde{n_{n}}(k)\right]^{T}$, where:
@@
${\widetilde{H}}=\left[\begin{array}[]{ccc}\widetilde{h_{11}}&\cdots&\widetilde{h_{1n}}\\ \vdots&\ddots&\vdots\\ \widetilde{h_{m1}}&\cdots&\widetilde{h_{mn}}\end{array}\right]$
@@
where $h_{ij}$ the action potential of the $j^{th}$ motor unit recorded at the channel $i$. The noise in model 1 includes both electronic interference and the activity of motor units at the skin surface, depicted by low-energy action potentials that cannot be separated.
The linear instantaneous model in Equation 1 is inverted to retrieve the matrix of extended and spatially whitened sources. The inversion of the matrix is performed with an iterative gradient descent optimisation procedure with a contrast function that maximises the independence, or sparseness [43], of the estimated sources. This method extracts the motor neuron discharge timings associated with individual motor neurons and the unique representation of the associated multi-channel action potentials. The estimated sources are trains of delta functions centered at motor neuron activation instants, with varying amplitudes due to the estimation process. To separate the discharge timings from the baseline noise, K-Means clustering is applied to the estimated sources, comparing each candidate peak with the surrounding peaks.
Several blind source separation approaches have been proposed in the literature [13, 44, 45, 46, 47, 48]. However, these methods often struggle to separate sources when their waveforms are similar, resulting in overlapping signals. Current contrast functions and non-linear metrics are not optimised to resolve these overlaps. We thus utilised a novel algorithm, Swarm-Contrastive Decomposition (SCD) [42], that addresses these limitations by using an adaptive contrast function that enhances spike sorting accuracy by adapting to the unique characteristics of spiking distributions.

### v2

@@ 4.1.1 Participants
Two healthy men, aged 39 and 30 years, were recruited for this study. Both participants had no history of neurological or musculoskeletal disorders that could affect motoneuron functionality. The small participant sample was due to the relatively complex implant procedure, which included three microelectrode arrays per participant, with ultrasound-guided insertion. All experimental procedures adhered to the ethical guidelines set by Imperial College London. The studies were performed according to the Declaration of Helsinki, with an informed consent form signed by all participants before each experiment (ICREC Project ID 19IC5640 for the intramuscular recordings; JRCO Project ID 18IC4685 for the MRI acquisition).
@@ 4.1.2 Experiments
HD-iEMG signals were recorded from forearm muscles using multi-channel intramuscular electrodes designed for acute recordings (13, 12) (Fig. 5A). Additionally, three surface electrode grids (IED 4mm) with 64 recording sites each (Fig. 5C) were positioned over the targeted muscles to concurrently acquire HD-sEMG signals. This configuration enabled concurrent measurement of the electrical activity both within the muscle and at the skin surface. The EMG signals were acquired using a multi-channel amplifier (OT-Bioelettronica, Torino, Italy). All EMG data were sampled at 10,240 Hz, and digitised with 16-bit resolution. The HD-EMG signals were bandpass filtered (Butterworth, order 5) between 100-4400 Hz (HD-iEMG) and 20-500Hz (HD-sEMG). Recordings were made in single-differential configuration with a reference electrode placed on the wrist.
The participants performed force-tracking tasks following trapezoidal force trajectories at 15 % of their MVC (Fig. 3B). These tasks were executed individually for each finger (thumb, index, middle, ring, and little finger) in flexion and extension. Additionally, forces were exerted by combinations of fingers (thumb and index; index and middle; middle and ring; index with middle and ring; thumb with index, middle, ring, and little finger) in flexion and extension, following the same trapezoidal profile. As shown in Figure 1, the fingers were constrained within the flexion and extension load cell, ensuring no dynamic movement between states was involved. During all tasks, participants maintained isometric contractions, focusing on the specified contractions. Each trial consisted of a 1-second initial rest period, followed by a 3-second ramp phase (5 %MVC/s), a 20-second constant-force hold, and a ramp-down phase (5 %MVC/s), concluding with a 1-second rest period. A 2-minute rest was provided between contractions to minimise fatigue. Two repetitions were recorded for each task.
@@
We classified the different tasks with a series of analyses. First, we attempted the classification of 16 classes that included all individual fingers in flexion and extension, and three combinations of two fingers (T-F, I-F, M-F, R-F, L-F, IM-F, MR-F, IMR-F, T-E, I-E, M-E, R-E, L-E, IM-E, MR-E, IMR-E). Then, we focused on the tasks mostly associated to the targeted muscles, which resulted in 12 classes: I-F, M-F, R-F, IM-F, MR-F, IMR-F, I-E, M-E, R-E, IM-E, MR-E, IMR-E. Finally, as a proof of concept of a specific application of the mapping system - the control of a prosthetic hand - we selected tasks that are functionally important for amputees and challenging to decode with state-of-the-art methods (15, 33). This group of tasks included power grip (TIMRL-F, TIMRL-E), pinch (TI-F, TI-E), and individual index/thumb finger activation (T-F, T-E, I-F, I-E), for a total of eight classes. All tasks were performed isometrically, with the corresponding fingers activated simultaneously. However, the power grip class was excluded for participant one due to the inability to correctly perform the coordinated contractions of the thumb, index, middle, ring, and little fingers. Specifically, participant one exhibited oscillations in the force applied by the little finger, with fluctuations of amplitude of around 7 %MVC. These force variations led to inconsistent activation of the motoneurons, which were subsequently removed by the automatic cleaning process. As a result, the last analysis resulted in six classes for participant one (T-F, I-F, TI-F, T-E, I-E, TI-E) and eight for participant two (T-F, I-F, TI-F, TIMRL-F, T-E, I-E, TI-E, TIMRL-E).
@@ 4.1.3 Instrumented platform
A finger force measurement platform was designed to be adaptable to different hand sizes, adjustable in wrist angle, and equipped with high-resolution and cost-effective load cells to capture force signals. A 10-mm polycarbonate board was used for the platform body because of its light weight. Precision laser-cutting of load cell rails was executed using CAD software. The platform features ten 10-kg TAL220 load cells (two per finger, one for flexion, and one for extension), selected based on the maximum observed male exertion forces of 7-8 kg (34, 35). The load cells are mounted on adjustable rails with rolling pins and connected to 3D-printed mounts to accommodate varying finger thicknesses. The EMG signals and the force data were acquired simultaneously by the Quattrocento amplifier. Force signals were amplified using Forza pre-amplifiers linked to the amplifier. A 3D-printed holder housed the 10 pre-amplifiers, optimising the setup for accurate force measurement across various hand sizes.
@@ 4.2 Classification
We examined the accuracy in classifying intended motor tasks into predefined categories by comparing two approaches: one that used decomposed motoneuron activity (i.e., the timings of motoneuron discharge events) and another that used global features derived from the interference signal (e.g., signal amplitude).
@@
To classify tasks using motoneuron behaviour, the HD-iEMG signals were decomposed into the contributing motoneurons. The filters associated with each source (motoneuron) were aggregated into a comprehensive dictionary of filters for each task. Figure 4 panel A provides a schematic representation of this filter bank. In the diagram, each colour corresponds to a different class label, indicating the specific task being represented. The width of each bar reflects the number of filters associated with that task, illustrating how many motoneuron filters contribute to the classification of each specific task. The zoomed-in section within the magenta box offers a closer look at a representative filter from the pooled filter bank. This filter visualises the amplitude of the motoneuron across different channels, providing insights into how the neural activity is distributed spatially across the electrode array. Conversely, for global EMG classification, the root mean square of the EMG signals served as the feature for a C-Support Vector Machine classifier (Fig. 4C). To mimic a real-life mapping system, the calibration phase was conducted using the first recorded repetition, and the testing phase was performed using a repetition recorded later in the day.
4.2.1 Finger task classification based on motoneuron behaviour
@@
Motor unit spike trains were extracted through blind source separation of HD-iEMG recordings obtained during isometric finger tasks at 15 %MVC (see Appendix A for more details on EMG decomposition). The tasks included individual finger flexion and extension as well as multi-finger combinations performed isometrically. The decomposition process produced task-specific sets of motoneuron filters, which were pooled to create a comprehensive bank of motoneuron action potential filters $\boldsymbol{\mathbb{F}}_{i}\in\mathbb{R}^{N\times M}$, where $N$ represents the total number of extracted motoneuron filters and $M$ the number of EMG channels. Each motoneuron filter $\boldsymbol{\mathbb{F}}_{i}$ in this set was labelled according to the specific finger movement task it was derived from (e.g., thumb flexion corresponding to class 0).
@@
These reference motoneuron filters from the 15 %MVC contraction were then applied to the multi-channel iEMG signals $\mathbf{E}j\in\mathbb{R}^{M\times D_{R}}$, where $D_{R}$ is the duration of the recording, recorded during a separate 15 % MVC trial involving all fingers performing both flexion and extension contractions. This produced the corresponding discharge timings, represented as IPTs $\mathbf{I}{ij}\in\mathbb{R}^{N\times D_{R}}$ for each motoneuron $i$ and task $j$. Each IPT $\mathbf{I}_{ij}$ was clustered to separate the peaks (i.e., the source) from the noise components. Only motoneurons with robust clustering quality (silhouette value >0.9), physiological discharge periodicity (coefficient of variation <35%), and firing rates within the physiological range for isometric contractions (mean discharge rate <30 Hz) were retained for further processing. The discharge timings of these units were temporally segmented into 100-ms time bins with 50% overlap. This window was selected to be long enough for reliable classification while remaining short enough to minimize delay in real-time control applications. The windowing resulted in a binary matrix $\mathbf{B}\in\mathbb{R}^{N\times K}$, where $K$ represents the number of bins. For each time bin $k$, the corresponding class label was assigned according to the filter $\boldsymbol{\mathbb{F}}_{i}$ that produced the source, creating a final label vector $\mathbf{L}\in\mathbb{R}^{K}$. If no predictions were made for a given bin (i.e., no motoneurons were active), it was labeled as N/A. Finally, for each time bin, the class with the majority of votes was assigned to that bin. If there were no predictions overall for a bin, it was labeled as N/A and counted as a misclassification in the final accuracy calculation. The method, therefore, relies on unique motoneurons being consistently active throughout the entire task. Figure 4A,B shows a schematic of the method.
@@ 4.2.2 Finger classification based on global EMG
For myoelectric decoding of dexterous hand tasks, a linear SVM classifier with an L2 regularisation penalty factor of 0.1 was employed (Fig. 4C). The SVM classifier is a widely used method in EMG pattern recognition (36, 37, 38). The HD-EMG signals were bandpass filtered (Butterworth, order 5) between 100-4400 Hz (HD-iEMG) and 20-500 Hz (HD-sEMG) to remove noise. Global normalisation factors based on the maximum absolute value across the pooled training set were derived and applied to the test set for consistent scaling. Outlier channels with baseline noise exceeding 3 standard deviations from the training mean were removed from all training and testing data. The pre-processed EMG was segmented into 100-ms intervals, with 50% overlap using a sliding window, preparing the bins similarly to how they were organised for the classification based on motoneurons. The RMS values of the EMG were extracted as the input features. This was done in the same way for both surface and intramuscular high-density recordings. Moreover, to simulate a lower-density iEMG electrode configuration, we selected the first and last channel for each electrode, resulting in an IED of 2 cm (Fig. 5B). The difference between the two electrodes was computed for each array, resulting in 3 feature channels. This configuration was chosen to emulate structures like the implantable myoelectric sensors (IMES) electrodes, currently being tested to control multifunction hand prostheses (39, 40, 41). For the surface electromyography (sEMG), we selected eight channels in the central column of the electrode grid, mimicking the configurations found in commercial low-density systems such as the MYOband (Fig. 5D).
@@ 5 Acknowledgments
A.G. was supported by UK Research and Innovation [UKRI Centre for Doctoral Training in AI for Healthcare grant number EP/S023283/1] and by Huawei Technologies Research & Development (UK) Limited. J.I.P. was supported by project ECHOES (ERC Starting Grant 101077693) and by a Consolidación Investigadora grant (CNS2022-135366) funded by MCIN/AEI/10.13039/ 501100011033 and UE’s NextGenerationEU/PRTR funds. S.M. was supported by Chalmers Life Science Engineering Area of Advance. A.K. was supported by NISNEM (EPSRC EP/T020970/1). D.F. was supported by NaturalBionicS (ERC Synergy 810346).
@@ Appendix A Decomposition
The EMG signals were decoded into the discharges of the corresponding motoneurons via a blind source separation algorithm (42).
Multi-channel EMG signals represent combined mixtures of motoneuron spike trains. Each recorded signal, i.e., each EMG channel, is the summation of the convolution of unknown finite impulse response filters, corresponding to motor unit action potentials, with their respective sources, i.e., the sequences of motoneuron discharge timings. To make the deconvolution (i.e., EMG decomposition) possible, this combined mixture can be converted into a linear instantaneous mixture by extending the sources to include the original $N$ sources and their delayed versions, with delays ranging from 1 to the filter length $L$. Similarly, the $M$ original observations are extended by a factor $R$ to ensure there are more observations than sources (43).
For each time sample $(k)$, indicating the sources, observations and noise, respectively, with $\underset{=}{s}(k)=\left[s_{1}(k),s_{2}(k),\ldots,s_{N}(k)\right]^{T},\underset{=}{x}(k)=\left[x_{1}(k),x_{2}(k),\ldots,x_{M}(k)\right]^{T}$ and $\underset{=}{n}$ $(k)=\left(n_{1}(k),n_{2}(k),\ldots,n_{N}(k)\right)^{T}$, where $T$ is the transpose operator, the extended model is as follows:
$\underset{=}{\widetilde{x}}(k)=[\underset{=}{\widetilde{H}}\widetilde{s}(k)+\underset{=}{\widetilde{n}}(k)]\quad\text{for}\quad k=1,\ldots,D_{R}$ (1)
where $D_{R}$ is the duration of the recording in samples, and with the extended sources, observations and noise as $\widetilde{s}(k)=\left[\widetilde{s_{1}}(k),\widetilde{s_{2}}(k),\ldots,\widetilde{s_{n}}(k)\right]^{T}$, $\widetilde{x}(k)=\left[\widetilde{x_{1}}(k),\widetilde{x_{2}}(k),\ldots,\widetilde{x_{M}}(k)\right]^{T}$ and $\widetilde{n}(k)=\left[\widetilde{n_{1}}(k),\widetilde{n_{2}}(k),\ldots,\widetilde{n_{N}}(k)\right]^{T}$, where:
@@
$\underset{=}{\widetilde{H}}=\left[\begin{array}[]{ccc}\widetilde{h_{11}}&\cdots&\widetilde{h_{1N}}\\ \vdots&\ddots&\vdots\\ \widetilde{h_{M1}}&\cdots&\widetilde{h_{MN}}\end{array}\right$]
@@
where $h_{ij}$ the action potential of the $j^{th}$ motoneuron recorded at the channel $i$. The noise in model 1 includes both electronic interference and the activity of motoneurons at the skin surface, depicted by low-energy action potentials that cannot be separated.
The linear instantaneous model in Equation 1 is inverted to retrieve the matrix of extended and spatially whitened sources. The inversion of the matrix is performed with an iterative gradient descent optimisation procedure with a contrast function that maximises the independence, or sparseness (43), of the estimated sources. This method extracts the motoneuron discharge timings associated with individual motoneurons and the unique representation of the associated multi-channel action potentials. The estimated sources are trains of delta functions centered at motoneuron activation instants, with varying amplitudes due to the estimation process. To separate the discharge timings from the baseline noise, K-Means clustering is applied to the estimated sources, comparing each candidate peak with the surrounding peaks.
Several blind source separation approaches have been proposed in the literature (13, 44, 45, 46, 47, 48). However, these methods often struggle to separate sources when their waveforms are similar, resulting in overlapping signals. Current contrast functions and non-linear metrics are not optimised to resolve these overlaps. We thus utilised a novel algorithm, Swarm-Contrastive Decomposition  (42, 49), that addresses these limitations by using an adaptive contrast function that enhances spike sorting accuracy by adapting to the unique characteristics of spiking distributions.